OpenStudy (anonymous):

Simplify 4^sqrt 400/4^sqrt 5. how to?

OpenStudy (baru):

can you use the equation button? its hard to interpret what you have typed

OpenStudy (daniel.ohearn1):

If you can take a screenshot of the problem that would help, attach it as a file too..

OpenStudy (jiteshmeghwal9):

\[\LARGE{4^{\sqrt{400}} \over {4^\sqrt{5}}}\]

OpenStudy (jiteshmeghwal9):

\[\LARGE{4^{20-\sqrt{5}}}\]as\[\LARGE{{a^b \over a^c}=a^{b-c}}\]

OpenStudy (denonakavro):

\[4^{400}\div4^{5} \] Just subtract 400-5 = 395 SO the answer will be 4^395

OpenStudy (faiqraees):

@denonakavro Please check the question again. The exponent with the second number is \[\sqrt{5}\] not simply 5

OpenStudy (whpalmer4):

@jiteshmegwhal9 has it correct. The step left out is reducing \(\sqrt{400}\) to \(20\) \[\frac{4^{\sqrt{400}}}{4^{\sqrt{5}}}=\frac{4^{20}}{4^{\sqrt{5}}}\]and because\[\frac{b^m}{b^n} = b^{m-n}\]we can then rewrite that as \[4^{20-\sqrt{5}}\] The approximate value of that is \(49539630091.22790143784586\), if you were wondering :-)